# Understanding Dynamic Scenes in Ego Centric 4D Point Clouds

### v2

@@ Abstract
{strip}[Uncaptioned image]
Figure 1: We introduce a novel QA benchmark EgoDynamic4D and an end-to-end spatio-temporal reasoning framework.
@@
1 Introduction
With the rapid advancement of embodied AI and human–machine interaction technologies, understanding dynamic 4D scenes from an egocentric viewpoint—integrating 3D spatial dimensions and the temporal dimension—has emerged as a pivotal research challenge. Unlike conventional third-person video analysis (Pang et al. 2021; Song et al. 2021; Pan et al. 2021; Li, Niu, and Zhang 2022), egocentric videos exhibit high dynamics, frequent scene changes, and rich interactive behaviors, requiring models not only to capture the wearer’s movement but also to perceive and reason about surrounding people, objects, and their evolving relationships (Li et al. 2021; Yu et al. 2023; Fan et al. 2017; Lee, Lee, and Choi 2023; Lin et al. 2022; Pramanick et al. 2023). Applications such as robotic perception, AR, and autonomous driving demand efficient and accurate ego-centric scene understanding to empower the next generation of embodied agents.
Although landmark egocentric datasets Ego4D (Grauman et al. 2022), EgoExo4D (Grauman et al. 2024), HD-Epic (Perrett et al. 2025), and 3D datasets (Dai et al. 2017; Wang et al. 2024a; Chen, Chang, and Nießner 2020) have driven progress in action recognition, object locating, and 3D scene analysis, they suffer from the following key limitations: (1) Incomplete 4D annotations and lack of dynamic content: Egocentric datasets often lack temporally aligned 3D bounding boxes and trajectories, while 3D datasets focus on static scenes without moving objects or agents—limiting the study of real-world dynamics. (2) Limited evaluation of temporal reasoning: Existing benchmarks emphasize short-term or moment-based tasks and lack protocols to assess reasoning over continuous object motion or interaction. (3) Incomplete multimodal evaluation: While some works explore temporal scene graphs (Yang et al. 2023), they focus on representation construction rather than end-to-end multimodal reasoning, and do not support QA-based evaluation of dynamic 4D scenes.
@@ 2 Related Work
3DLLM (Hong et al. 2023) and 3UR-LLM (Xiong et al. 2025) take point clouds as inputs for tasks like description and localization. LLaVA3D (Zhu et al. 2024) and Video3DLLM (Zheng, Huang, and Wang 2025) introduce ”3D patches” by embedding 2D CLIP features (Radford et al. 2021) with 3D positional encodings. GPT4Scene (Qi et al. 2025) uses BEV imagery (Liu et al. 2022; Li et al. 2022a) and consistent object IDs to reason about spatial relations. LSceneLLM (Zhi et al. 2025) employs LLM attention to identify relevant regions for fine details. Chat-3D (Wang et al. 2023) and Chat-Scene (Huang et al. 2024) use object identifiers to enable interactive understanding. While effective on static 3D scenes, these approaches lack mechanisms to capture temporal dynamics for 4D reasoning.
@@
CoT has been extended to reasoning in vision. Video‑CoT(Wang et al. 2024b) introduces QA pairs with CoT to benchmark reasoning over temporal video content. SpatialCoT(Liu et al. 2025) targets embodied spatial reasoning by aligning vision–language inputs with spatial coordinates. (Zeng et al. 2025) extends this idea to autonomous driving. Likewise, EgoDynamic4D provides rich CoT data for spatio‑temporal reasoning.
@@ 3.1 Overview
The EgoDynamic4D benchmark in Figure 1 interprets complex scene dynamics, track moving objects, and analyze interactions from a first-person perspective. Given an input 4D egocentric scene $S=\{F_{t}\}_{t=1}^{T}$, where $F_{t}=(I_{t},D_{t})$ represents a frame at timestamp $t$ with image $I_{t}$, depth map $D_{t}$, and a natural language question $Q$ about the scene, the model predicts answer $A$. The answer $A$ can be a categorical label, numerical value, or descriptive text, depending on the specific type of question. We organize the QA tasks into three domains: Scene Descriptions capture semantics of the environment. This includes object-captioning, which requires understanding of objects and their surroundings given 2D/3D bounding boxes at specific time. Momentary Dynamics focus on real-time spatial relations and short-term motion cues. This domain includes object-centric tasks such as dynamic-scene, relative-position, and current-object-property, as well as agent-centric tasks such as agent-velocity and multi-agent-relation. Durative Dynamics address longer-term changes and interactions across time. Object-centric tasks include temporary-static-objects, most-active-object, and motion-sequence, while agent-centric tasks include agent-trajectory, agent-grab-object, and agent-motion-status. Detailed definitions are shown in the appendix.
@@ 3.3 Data Distribution
Figure 5: The end-to-end framework which encode dynamic 4D scenes based on the ego-centric videos.
@@ 4.1 Instance and Timestamps Enhanced Point-level Feature Extraction
obj.
caption
(bleu4)↑
dyn.
scene
(f1)↑
rel.
pos.
(acc%)↑
curr.
obj.
prop.
(acc%)↑
agent
vel.
(acc%)↑
multi
agent
rel.
(acc%)↑
motion
seq.
(acc%)↑
most
active
obj.
(acc%)↑
temp.
static
obj.
(f1)↑
agent
traj.
(acc%)↑
agent
motion
status
(acc%)↑
agent
grab
obj.
(acc%)↑
@@ 4.2 Feature Fusion
A learnable matrix $W_{ins}\in\mathbb{R}^{d_{vis}\times d_{ins}}$ projects instance embedding $\overline{f_{vis,v}}$ into hidden dimension, then it is fused with time embeddings and original position embeddings via self-attention, adding to visual feature together.
@@
obj.
caption
(bleu4)↑
dyn.
scene
(f1)↑
rel.
pos.
(acc%)↑
curr.
obj.
prop.
(acc%)↑
agent
vel.
(acc%)↑
multi
agent
rel.
(acc%)↑
motion
seq.
(acc%)↑
most
active
obj.
(acc%)↑
temp.
static
obj.
(f1)↑
agent
traj.
(acc%)↑
agent
motion
status
(acc%)↑
agent
grab
obj.
(acc%)↑
@@
We compress the sequence of camera poses $(x,y,z,q_{x},q_{y},q_{z},q_{w})\in\mathbb{R}^{7}$ into a compact embedding using a learnable attention mechanism. A sequence of $T$ poses are first project into hidden space $f_{cam}\in\mathbb{R}^{T\times d_{vis}}$, and then $M$ learnable query tokens $Q_{cam}\in\mathbb{R}^{M\times d_{vis}}$ attends to this sequence through cross-attention:
@@ 5.1 Implementation Details
We conduct experiments EgoDynamic4D (80% train, 20% test). Our model builds upon LLaVA3D (CLIP + LLaMA) (Zhu et al. 2024) with frozen backbones, unfreezing only the proposed modules ($d_{ins}$=8, MM=8) and LoRA parameters (rank=8, alpha=16). We set sample fps to 5, and train using AdamW (learning rate 5e-5) for 2 epochs on 8 × RTX 4090 (24GB) with batch size 1 per GPU.
@@ 5.2 Results on EgoDynamic4D
We evaluate our method on the EgoDynamic4D dataset (ADT and THUD++ subsets), and benchmark it against representative baselines. Since 4D LLMs like LLaVA-4D (Zhou and Lee 2025) are not publicly available yet, we evaluated 3D LLMs including LLaVA-3D (Zhu et al. 2024), Video3DLLM (Zheng, Huang, and Wang 2025), LL3DA (Chen et al. 2024), Chat-Scene (Huang et al. 2024), 3DLLM (Hong et al. 2023) and VG-LLM (Zheng et al. 2025). To ensure a fair comparison under limited resources, we fine-tune all LLM models with the same LoRA configuration, while keeping other components in original implementation settings. Evaluation follows the benchmark’s protocol (details in the appendix), with thresholds settings: speed error within 0.05m/s0.05~m/s, direction error within 0.5rad0.5~rad, IOU above $0.1$, distance and position errors within 0.1m0.1~m.

### v3

@@ Abstract
1 Introduction
With the rapid advancement of embodied AI and human–machine interaction technologies, understanding dynamic 4D scenes from an egocentric viewpoint—integrating 3D spatial dimensions and the temporal dimension—has emerged as a pivotal research challenge. Unlike conventional third-person video analysis (Pang et al. 2021; Song et al. 2021; Pan et al. 2021; Li et al. 2022a), egocentric videos exhibit high dynamics, frequent scene changes, and rich interactive behaviors, requiring models not only to capture the wearer’s movement but also to perceive and reason about surrounding people, objects, and their evolving relationships (Li et al. 2021; Yu et al. 2023; Fan et al. 2017; Lee et al. 2023; Lin et al. 2022; Pramanick et al. 2023). Applications such as robotic perception, AR, and autonomous driving demand efficient and accurate egocentric scene understanding to empower the next generation of embodied agents.
@@ 1 Introduction
Although landmark egocentric datasets Ego4D (Grauman et al. 2022), EgoExo4D (Grauman et al. 2024), HD-Epic (Perrett et al. 2025), and 3D datasets (Dai et al. 2017; Wang et al. 2024a; Chen et al. 2020) have driven progress in action recognition, object locating, and 3D scene analysis, they suffer from the following key limitations: (1) Incomplete 4D annotations and lack of dynamic content: Egocentric datasets often lack temporally aligned 3D bounding boxes and trajectories, while 3D datasets focus on static scenes without moving objects or agents—limiting the study of real-world dynamics. (2) Limited evaluation of temporal reasoning: Existing benchmarks emphasize short-term or moment-based tasks and lack protocols to assess reasoning over continuous object motion or interaction. (3) Incomplete multimodal evaluation: While some works explore temporal scene graphs (Yang et al. 2023), they focus on representation construction rather than end-to-end multimodal reasoning, and do not support QA-based evaluation of dynamic 4D scenes.
@@ 2 Related Work
3DLLM (Hong et al. 2023) and 3UR-LLM (Xiong et al. 2025) take point clouds as inputs for tasks like description and localization. LLaVA3D (Zhu et al. 2024) and Video3DLLM (Zheng et al. 2025b) introduce 3D patches by embedding 2D CLIP features (Radford et al. 2021) with 3D positional encoding. GPT4Scene (Qi et al. 2025) uses BEV images and consistent object IDs to reason about spatial relations. LSceneLLM (Zhi et al. 2025) employs LLM attention to identify relevant regions for fine details. Chat-3D (Wang et al. 2023) and Chat-Scene (Huang et al. 2024) use object identifiers to enable interactive understanding. While effective on static 3D scenes, these approaches lack mechanisms to capture temporal dynamics for 4D reasoning.
@@
CoT has been extended to reasoning in vision, like (Yuan et al. 2023) and  (Yuan et al. 2024). Video‑CoT (Wang et al. 2024b) introduces QA pairs with CoT to benchmark reasoning over temporal video content. SpatialCoT (Liu et al. 2025) targets embodied spatial reasoning by aligning vision–language inputs with spatial coordinates. (Zeng et al. 2025) extends this idea to autonomous driving. Likewise, EgoDynamic4D provides rich CoT data for spatio‑temporal reasoning.
@@ 3.1 Overview
The EgoDynamic4D benchmark in Figure LABEL:fig:dataset-overview interprets complex scene dynamics, track moving objects, and analyze interactions from a first-person perspective. Given an input 4D egocentric scene $S=\{F_{t}\}_{t=1}^{T}$, where $F_{t}=(I_{t},D_{t})$ represents a frame at timestamp $t$ with image $I_{t}$, depth map $D_{t}$, and a natural language question $Q$ about the scene, the model predicts answer $A$. The answer $A$ can be a categorical label, numerical value, or descriptive text, depending on the specific type of question. We organize the QA tasks into three domains: Scene Descriptions capture semantics of the environment. This includes object-captioning, which requires understanding of objects and their surroundings given 2D/3D bounding boxes at specific time. Momentary Dynamics focus on real-time spatial relations and short-term motion cues. This domain includes object-centric tasks such as dynamic-scene, relative-position, and current-object-property, as well as agent-centric tasks such as agent-velocity and multi-agent-relation. Durative Dynamics address longer-term changes and interactions across time. Object-centric tasks include temporary-static-objects, most-active-object, and motion-sequence, while agent-centric tasks include agent-trajectory, agent-grab-object, and agent-motion-status. Detailed definitions are shown in the appendix.
@@ 3.3 Data Distribution
Figure 5: The end-to-end framework which encode dynamic 4D scenes based on the egocentric videos.
@@ 4.1 Instance and Timestamps Enhanced Point-level Feature Extraction
obj. caption (bleu4)↑ dyn. scene (f1)↑ rel. pos. (acc%)↑ curr. obj. prop. (acc%)↑ agent vel. (acc%)↑ multi agent rel. (acc%)↑ motion seq. (acc%)↑ most active obj. (acc%)↑ temp. static obj. (f1)↑ agent traj. (acc%)↑ agent motion status (acc%)↑ agent grab obj. (acc%)↑
@@ 4.2 Feature Fusion
A learnable matrix $W_{ins}\in\mathbb{R}^{d_{vis}\times d_{ins}}$ projects instance embedding $\overline{f_{vis,v}}$ into hidden dimension, then it is fused with time embeddings and original position embeddings via self-attention (SA), adding to visual feature together.
@@
obj. caption (bleu4)↑ dyn. scene (f1)↑ rel. pos. (acc%)↑ curr. obj. prop. (acc%)↑ agent vel. (acc%)↑ multi agent rel. (acc%)↑ motion seq. (acc%)↑ most active obj. (acc%)↑ temp. static obj. (f1)↑ agent traj. (acc%)↑ agent motion status (acc%)↑ agent grab obj. (acc%)↑
@@
We compress the sequence of camera poses $(x,y,z,q_{x},q_{y},q_{z},q_{w})\in\mathbb{R}^{7}$ into a compact embedding using a learnable attention mechanism. A sequence of $T$ poses are first project into hidden space $f_{cam}\in\mathbb{R}^{T\times d_{vis}}$, and then $M$ learnable query tokens $Q_{cam}\in\mathbb{R}^{M\times d_{vis}}$ attends to $f_{cam}$ through cross-attention (CA):
@@ 5.1 Implementation Details
We conduct experiments on EgoDynamic4D (80% train, 20% test). Our model builds upon LLaVA3D (CLIP + LLaMA) (Zhu et al. 2024) with frozen backbones, unfreezing only the proposed modules ($d_{ins}$=8, MM=8) and LoRA parameters (rank=8, alpha=16). We set sample fps to 5, and train using AdamW (learning rate 5e-5) for 2 epochs on 8 × RTX 4090 (24GB) with batch size 1 per GPU.
@@ 5.2 Results on EgoDynamic4D
We evaluate our method on the EgoDynamic4D dataset (ADT and THUD++ subsets), and benchmark it against representative baselines. Since 4D LLMs like LLaVA-4D (Zhou and Lee 2025) are not publicly available yet, we evaluated 3D LLMs including LLaVA-3D (Zhu et al. 2024), Video3DLLM (Zheng et al. 2025b), LL3DA (Chen et al. 2024), Chat-Scene (Huang et al. 2024), 3DLLM (Hong et al. 2023) and VG-LLM (Zheng et al. 2025a). To ensure a fair comparison under limited resources, we fine-tune all LLM models with the same LoRA configuration, while keeping other components in original implementation settings. Evaluation follows the benchmark’s protocol (details in the appendix), with thresholds settings: speed error within 0.05m/s0.05~m/s, direction error within 0.5rad0.5~rad, IOU above $0.1$, distance and position errors within 0.1m0.1~m.
@@ 6 Discussion and Conclusion
Acknowledgments
This work was sponsored by the National Key R&D Program of China (No. 2024YFF0907803), the National Natural Science Foundation of China (No. 62576308), and the Fundamental Research Funds for the Central Universities (No. 226-2025-00167).